OpenStudy (anonymous):

(x-5)/(x+1) = x-5)/(x+3)

OpenStudy (anonymous):

\[\frac{x-5}{x+1}=\frac{x-5}{x+3}\Rightarrow x=5,\ as\ x+1 \neq x+3\]